OpenStudy (anonymous):

Find (a) the elasticity of demand and (b) the range of prices for which the demand is elastic (E , -1). f(p) = 200(30 - p)

OpenStudy (anonymous):

@abb0t an u help?

OpenStudy (abb0t):

Is this econ?

OpenStudy (anonymous):

calculus my teacher said to use this E = p ⁄ f(p) * f'(p) and f(p) = 200(30 - p).

OpenStudy (abb0t):

So you're being asked to find the derivative (change in elasticity) right?

OpenStudy (anonymous):

this is what i did before and she said to correct it p = 10 f(10) = 4000 p = f(20) = 2000 then the elasticity is 4000 – 2000 over 4000 + 2000/ 10 – 20 over 10 + 20 = 1/3 over -1/3 So the elasticity is -1

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

Can you rewrite the formula using latex, I see you used some type of chain rule or product rule there. Because if you take the derivative of what I see, all you get is 6000-200 = 5800 and you can't evaluate a constant from E <p < -1

OpenStudy (anonymous):

okay and how would i rewrite it?

OpenStudy (abb0t):

use the equation editor option on the bottom of the text box

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(p) = 200(30 - p) \]

OpenStudy (anonymous):

\[ E = p ⁄ f(p) * f'(p) and f(p) = 200(30 - p)\]

OpenStudy (abb0t):

hmmm..strange question. LOL. It's not the math that im not understanding, it's the application of is. But I think that you're being asked to find "E" in which case you have: \[E = \frac{ p }{ (6000-200p) \times 5800 }\]

OpenStudy (anonymous):

okay hold on let me show you a video of it

OpenStudy (abb0t):

The 5800 is simply the derivative of the function, f(p)

OpenStudy (abb0t):

OHHHHHHHHHH I see what they did there. That's a good video, very similar to your question. What are you having trouble understanding?

OpenStudy (anonymous):

How to use that formula my teacher game me and also I'm not seeing how i did it wrong from the beginning?

OpenStudy (abb0t):

it says that revenue is \( p \times f(p)\) whcih in your case is \(p (6000-200p) = 6000p-200p^2\)

OpenStudy (abb0t):

take the derivative of that.

OpenStudy (anonymous):

the derivative of all of 600p - 200p^2

OpenStudy (anonymous):

i get 600 - 400p

OpenStudy (abb0t):

@Mertsj

OpenStudy (anonymous):

oh my bad its 6000

OpenStudy (anonymous):

i get -400(p - 15)